# Benchmarking Spatiotemporal Reasoning in LLMs and Reasoning Models: Capabilities and Challenges

### v2

@@ 1 Introduction
Cyber-Physical Systems (CPS) play a critical role in numerous daily and industrial applications, enabling applications like precision agriculture, environmental monitoring, and smart city infrastructure. Effective operation in these domains relies on spatiotemporal intelligence: the ability to perceive dynamic environments, understand spatiotemporal relationships, and integrate this understanding with world knowledge to make informed decisions. For instance, mobile robots need to fuse sensor data (e.g., range sensor) to navigate cluttered spaces. For various CPS applications, the common denominator is the necessity for reasoning jointly over space, time, and world knowledge.
Recent progress in large language models (LLMs) and large reasoning models (LRMs) openai2025o3o4mini ; jaech2024openai ; guo2025deepseek suggests their potential as general-purpose agents for CPS applications. Yet the community lacks rigorous benchmarks that consider the complexity in CPS applications. As shown in Figure 1, spatiotemporal reasoning in CPS typically proceeds through three stages: (1) First, state estimation. The system must perceive its environment and understand the current situation. This step involves determining what is happening, where, and when. (2) Reasoning over estimated states. Once a set of current and past states is estimated, the system needs to understand how these states interrelate across space and time. (3) Finally, world-knowledge-aware reasoning. For complex real-world scenarios, the estimated states and their spatiotemporal relationships are often interpreted with contextual information and world knowledge to make more nuanced inferences, predictions, or decisions.
However, existing benchmarks li2024stbench ; zhan2025open3dvqa ; mooney2023towards fall short of capturing this comprehensive, structured view of spatiotemporal reasoning. To address this gap, we introduce STARK (SpatioTemporal reAsoning benchmaRK), a hierarchical benchmark that measures how well models can plan and execute spatiotemporal reasoning pipelines. We designed STARK with the following features:
  • Task novelty. STARK includes under-explored problems such as field-variable prediction, spatiotemporal localization, and tracking that intertwine spatial geometry and temporal operation.
@@
  • Task difficulty. Unlike prior multiple-choice suites li2024stbench , STARK includes open-ended challenges: the model must output a numeric estimate given the reasoning challenges instead of picking from canned answers. Specifically, we evaluate the case where the model directly answers (DA) or invokes a Python code interpreter (CI) supplied at inference time. The latter serves as a proxy for evaluating the model’s capability of using external tools, an essential indicator of its proficiency as an AI agent. This design effectively exposes both the model’s direct reasoning ability and its competency in tool usage.
@@
  • Open resources. All data [^1] https://huggingface.co/datasets/prquan/STARK_10k and evaluation code [^2] https://github.com/nesl/STARK_Benchmark/ are publicly available to facilitate reproducibility and foster more capable spatiotemporal reasoning systems. By exposing where LLMs and LRMs excel and fall short, STARK can inspire more effective solutions for intelligent CPS applications.
@@ 2 Related works
GeoQAxu2024evaluating Coordinate $\times$ $\times$ $\times$ $\times$ DA
GeoQAmooney2023towards Coordinate $\times$ $\times$ $\times$ $\times$ $\times$ DA
TemporalQAtan2023towards Text $\times$ $\times$ $\times$ $\times$ $\times$ DA
Open3DVQAzhan2025open3dvqa Image $\times$ $\times$ $\times$ $\times$ DA
STBenchli2024stbench Coordinate $\times$ $\times$ $\times$ $\times$ DA
@@
Spatial, temporal, and spatiotemporal benchmark. GeoQA xu2024evaluating ; mooney2023towards evaluate symbolic (e.g., north-of) and geometry reasoning but ignore interaction with different sensor modalities and temporal reasoning. Temporal-QA tan2023towards ; wei2023menatqa focus on event ordering and interval logic in pure text, yet they do not include spatial constraints or open-ended numeric estimation. STBenchli2024stbench combines both spatial and temporal axes but either collapses tasks into multiple-choice or ignores other common spatial estimation tasks, such as field variable prediction, localization, or tracking. In addition, it lacks the evaluation of a spatiotemporal logic framework (Tier 2 in STARK) and reasoning with both algorithmic computation and world knowledge (Tier 3 in STARK). Vision-LLMs cheng2025v ; zhan2025open3dvqa are evaluated on spatiotemporal reasoning tasks using egocentric videos or images, focusing on object recognition, temporal event ordering, and coarse spatial grounding. However, it focuses on a single deployed sensor and lacks explicit geometric computation and multi-sensor reasoning. Lastly, none of the above works attempt to evaluate LRMs or the effectiveness of using CI.
Spatial and temporal logic frameworks. Spatial and temporal relations in STARK build directly on two classical frameworks that formally describe spatial and temporal operations. For space, we adopt the dimensionally-extended nine-intersection model (DE-9IM) clementini1993small ; clementini1994modelling implemented in ArcGIS scott2009spatial ; johnston2001using , whose topological predicates exhaustively categorize how any two geometries relate and form the backbone of modern GIS query engines, including intersects, contains, within, touches, overlaps, and their complements. For time, we rely on Allen’s interval algebra allen1983maintaining , whose 13 atomic relations provide a mutually exclusive, jointly exhaustive vocabulary for reasoning over intervals on a timeline (before, meets, overlaps, during, starts, finishes, equals and their converses). Together these frameworks constitute the symbolic operations for spatiotemporal inference. To our knowledge, no prior benchmark has directly evaluated whether LLMs can reason over both ArcGIS spatial predicates and Allen’s interval algebra in a structured setting. STARK fills this gap by incorporating both formalisms into open-ended, sensor-grounded tasks where models must apply formal spatial and temporal logic to realistic trajectories and interval structures.
@@ 3.1 Sensor modalities
To support realistic and diverse spatiotemporal tasks in STARK, we simulate five representative sensor modalities in Figure 2: Range (distance measurement, enabling multilateration), Bearing (angular measurement, supporting triangulation), Range & Bearing (jointly output both distance and angle), Proximity (binary output for coarse localization via set intersection), and Event-based (timestamped detections for spatiotemporal inference using time-of-arrival differences). These modalities reflect common real-world sensing systems and elicit a range of geometric and combinatorial reasoning strategies. Sensor readings are injected with modality-specific noise (e.g., 1% for range sensors to mimic LiDAR accuracy; full specifications in Appendix 7.4).
@@ 3.2.1 State estimation
Field variable prediction tasks aim to estimate sensor measurements across space and time, including spatial imputation (predicting a sensor’s value from neighboring sensors at the same time), temporal imputation (using historical and future values), spatiotemporal imputation (leveraging neighboring sensors’ histories), and spatiotemporal forecasting (predicting future values for neighboring sensors). We use real-world air quality data from PurpleAir noauthor_real-time_nodate and traffic data from CalTrans PeMS california_pems_nodate , and also synthesize temperature readings with controlled spatial properties.
@@
Spatial tracking addresses dynamic object localization along continuous trajectories. In our established four-sensor, $10\times 10$ plane simulation, the object movement produces a time-stamped trajectory and sensor measurement. Inspired by SCAAT tracking welch1997scaat , we introduce partial observability via random sensor dropout. This forces the model to handle incomplete, noisy data (e.g., only two active sensors) by historical trajectory patterns and motion continuity to infer position, reflecting real-world challenges in autonomous navigation. Appendix 7.3 details trajectory simulation.
Temporal tracking estimates the timing of events occurring along a moving trajectory. We adopt the same prior sensor and environment configuration similar to SCAAT welch1997scaat .
@@ 3.2.2 Reasoning over states
Spatial relationship reasoning evaluates a model’s ability to infer qualitative spatial relationships between geometric objects. Spatial predicates defined in ArcGIS johnston2001using ; scott2009spatial lead to 35 possible combinations of geometry pairs and spatial relations (Appendix 7.5). For each instance, a pair of geometric objects is sampled and a specific relationship is generated using Shapelyshapely2007 .
Temporal interval reasoning tests a model’s ability with Allen’s interval algebra allen1983maintaining . We generate tasks by simulating two random intervals, labeling their temporal relationship, and creating a balanced negative counterpart.
@@ 3.2.3 Knowledge-aware reasoning
Using ArcGIS johnston2001using ; scott2009spatial and OpenStreetMap haklay2008openstreetmap , we produce the following challenges that require world knowledge consideration: (1) Landmark direction evaluates a model’s ability to infer directional relationships between real-world landmarks. (2) Landmark proximity tests a model’s spatial reasoning in urban contexts by reasoning the presence of amenities near landmarks (e.g., Is there a police station within 500 meters of Central Park?). (3) Route planning tests whether models can connect textual travel descriptions to real-world spatial layouts. The model must decide if a given route passes by a specified landmark. (4) For ETA calculation, the model is asked to determine, given a departure time and average speed, if arrival before a specified deadline is possible based only on textual descriptions. (5) Route segment duration test the model on whether a user could plausibly be present along the segment during a given time window.
Additionally, we resort to simulating user trajectories for controlled ground-truth user intent to address data scarcity and privacy concerns in real-world mobility datasets. (6) POI prediction simulates human mobility using a finite state machine (FSM)fernandez2012human ; kerr2011activity , where each state represents a point of interest (POI). We generate a synthetic seven-day trajectory at hourly intervals. To enhance realism, we integrate contextual factors, such as weather, time, traffic, social events, etc., which are also expressed in natural language, into the models. The model must predict the next likely POI, requiring reasoning over spatiotemporal traces and language context (Figure 3 (b)). (7) Similarly, intent prediction extends POI prediction by requiring the model to infer the user’s underlying intent behind each visit. The model must predicts activity category (e.g., going home, watching a movie).
@@ 3.3 Baseline generation
We establish first-principle-based baselines for each task using theoretically grounded solutions for localization, tracking, and field variable prediction (details in Appendix 7.6). We do not construct first-principle baselines for the remaining tasks, as using spatial libraries (e.g., Shapely shapely2007 ) and geospatial services (e.g., ArcGIS johnston2001using ) yields perfect accuracy. Nonetheless, the random-guess baseline for the tasks yields 0.5 accuracy due to a balanced distribution of positive and negative pairs.
@@ 4.1 Experiment Setup
We conduct evaluations on a subset comprising 1,315 questions in STARK, due to financial constraints. For all tasks except field variable prediction, we report the Root Mean Squared Error (RMSE) against the ground truth (for both open-ended calculation and true or false questions). For field variable prediction, where outputs vary in scale, we instead report the Root Mean Squared Percentage Error (RMSPE), because of differences in data magnitude across datasets. To ensure robustness to extreme outliers, we compute all metrics using the trimmed values unless stated otherwise, i.e., excluding the top and bottom 10% of the distribution.
For evaluating the models using a code interpreter (CI), we execute the code in a sandbox environment and report the relative change in error compared to direct answering (DA). This is calculated as the percentage change in RMSE and RMSPE due to using CI (lower values indicate better performance). Code execution was supported using scipy, numpy, shapely, and filterpy libraries 2020SciPy-NMeth ; harris2020array ; shapely2007 ; filterpy . All system prompts are provided in Table 5. Costs are shown in Table 6.

### v3

@@ 1 Introduction
Cyber-Physical Systems (CPS) play a critical role in numerous daily and industrial applications, enabling applications such as precision agriculture, environmental monitoring, and smart city infrastructure. Effective operation in these domains relies on spatiotemporal intelligence: the ability to perceive dynamic environments, understand spatiotemporal relationships, and integrate this understanding with world knowledge to make informed decisions. For instance, mobile robots need to fuse sensor data (e.g., range sensor) to navigate cluttered spaces. For various CPS applications, the common denominator is the necessity for reasoning jointly over space, time, and world knowledge.
Recent progress in large language models (LLMs) and large reasoning models (LRMs) OpenAI (2025); Jaech et al. (2024); Guo et al. (2025) suggests their potential as general-purpose agents for CPS applications. Yet the community lacks rigorous benchmarks that consider the complexity in CPS applications. As shown in Figure 1, spatiotemporal reasoning in CPS typically proceeds through three stages: (1) First, state estimation. The system must perceive its environment and understand the current situation. This step involves determining what is happening, where, and when. (2) Reasoning over estimated states. Once a set of current and past states is estimated, the system needs to understand how these states interrelate across space and time. (3) Finally, world-knowledge-aware reasoning. For complex real-world scenarios, the estimated states and their spatiotemporal relationships are often interpreted with contextual information and world knowledge to make more nuanced inferences, predictions, or decisions.
However, existing benchmarks Li et al. (2024); Zhan et al. (2025); Mooney et al. (2023) fall short of capturing this comprehensive, structured view of spatiotemporal reasoning. To address this gap, we introduce STARK (SpatioTemporal reAsoning benchmaRK), a hierarchical benchmark that measures how well models can plan and execute spatiotemporal reasoning pipelines. Besides, in addition to direct answering, our benchmark is specifically designed to assess tool usage competency, viewing it as an essential move to benchmark real-world CPS agents. We designed STARK with the following features:
  • Task novelty. STARK includes under‑explored problems such as field‑variable prediction, spatiotemporal localization, and tracking that intertwine spatial geometry and temporal operation.
@@
  • Task difficulty. Unlike prior multiple‑choice suites Li et al. (2024), STARK includes open‑ended challenges: the model must output a numeric estimate given the reasoning challenges instead of picking from canned answers. Specifically, we evaluate the case where the model directly answers (DA) or invokes a Python code interpreter (CI) supplied at inference time. The CI mode serves as a primary metric for evaluating the model’s capability to use external, deterministic tools, a core requirement for reliable CPS agents. This design effectively exposes both the model’s spatiotemporal reasoning ability and its competency in tool usage.
@@
  • Open resources. The benchmarks, STARK-L (14k samples for comprehensive evaluation  [^1] https://huggingface.co/datasets/prquan/STARK_10k) and STARK-S (1.3k samples for rapid usage [^2] https://huggingface.co/datasets/prquan/STARK_1k), and code [^3] https://github.com/nesl/STARK_Benchmark/ are publicly available to facilitate reproducibility and foster more capable spatiotemporal reasoning systems. By exposing where LLMs and LRMs excel and fall short, STARK can inspire more effective solutions for intelligent CPS applications.
@@ 2 Related works
GeoQAXu et al. (2024) Coordinate $\times$ $\times$ $\times$ $\times$ DA
GeoQAMooney et al. (2023) Coordinate $\times$ $\times$ $\times$ $\times$ $\times$ DA
TemporalQATan et al. (2023) Text $\times$ $\times$ $\times$ $\times$ $\times$ DA
Open3DVQAZhan et al. (2025) Image $\times$ $\times$ $\times$ $\times$ DA
STBenchLi et al. (2024) Coordinate $\times$ $\times$ $\times$ $\times$ DA
@@
Spatial, temporal, and spatiotemporal benchmark. GeoQA Xu et al. (2024); Mooney et al. (2023) evaluates symbolic (e.g., north‑of) and geometry reasoning but ignores interaction with different sensor modalities and temporal reasoning. Temporal-QA Tan et al. (2023); Wei et al. (2023) focus on event ordering and interval logic in pure text, yet they do not include spatial constraints or open-ended numeric estimation. STBenchLi et al. (2024) combines both spatial and temporal axes but either collapses tasks into multiple‑choice or ignores other common spatial estimation tasks, such as field variable prediction, localization, or tracking. In addition, it lacks the evaluation of a spatiotemporal logic framework (Tier 2 in STARK) and reasoning with both algorithmic computation and world knowledge (Tier 3 in STARK). Vision-LLMs Cheng et al. (2025); Zhan et al. (2025) are evaluated on spatiotemporal reasoning tasks using egocentric videos or images, focusing on object recognition, temporal event ordering, and coarse spatial grounding. However, it focuses on a single deployed sensor and lacks explicit geometric computation and multi-sensor reasoning. Lastly, none of the above works attempt to evaluate LRMs or the effectiveness of using CI.
Spatial and temporal logic frameworks. Spatial and temporal relations in STARK build directly on two classical frameworks that formally describe spatial and temporal operations. For space, we adopt the dimensionally‑extended nine‑intersection model (DE‑9IM) Clementini et al. (1993, 1994) implemented in ArcGIS Scott and Janikas (2009); Johnston et al. (2001), whose topological predicates exhaustively categorize how any two geometries relate and form the backbone of modern GIS query engines, including intersects, contains, within, touches, overlaps, and their complements. For time, we rely on Allen’s interval algebra Allen (1983), whose 13 atomic relations provide a mutually exclusive, jointly exhaustive vocabulary for reasoning over intervals on a timeline (before, meets, overlaps, during, starts, finishes, equals and their converses). Together these frameworks constitute the symbolic operations for spatiotemporal inference. To our knowledge, no prior benchmark has directly evaluated whether LLMs can reason over both ArcGIS spatial predicates and Allen’s interval algebra in a structured setting. STARK fills this gap by incorporating both formalisms into open-ended, sensor-grounded tasks where models must apply formal spatial and temporal logic to realistic trajectories and interval structures.
@@ 3.1 Sensor modalities
To support realistic and diverse spatiotemporal tasks in STARK, we simulate five representative sensor modalities in Figure 2: Range (distance measurement, enabling multilateration), Bearing (angular measurement, supporting triangulation), Range & Bearing (jointly output both distance and angle), Proximity (binary output for coarse localization via set intersection), and Event-based (timestamped detections for spatiotemporal inference using time-of-arrival differences). These modalities reflect common real-world sensing systems and elicit a range of geometric and combinatorial reasoning strategies. Sensor readings are injected with modality-specific noise (e.g., 1% for range sensors to mimic LiDAR accuracy; full specifications in Appendix A.4).
@@ 3.2.1 State estimation
Field variable prediction tasks aim to estimate sensor measurements across space and time, including spatial imputation (predicting a sensor’s value from neighboring sensors at the same time), temporal imputation (using historical and future values), spatiotemporal imputation (leveraging neighboring sensors’ histories), and spatiotemporal forecasting (predicting future values for neighboring sensors). We use real-world air quality data from PurpleAir 19 and traffic data from CalTrans PeMS State of California , and also synthesize temperature readings with controlled spatial properties.
@@
Spatial tracking addresses dynamic object localization along continuous trajectories. In our established four-sensor, $10\times 10$ plane simulation, the object movement produces a time-stamped trajectory and sensor measurement. Inspired by SCAAT tracking Welch and Bishop (1997), we introduce partial observability via random sensor dropout. This forces the model to handle incomplete, noisy data (e.g., only two active sensors) by historical trajectory patterns and motion continuity to infer position, reflecting real-world challenges in autonomous navigation. Appendix A.3 details trajectory simulation.
Temporal tracking estimates the timing of events occurring along a moving trajectory. We adopt the same prior sensor and environment configuration similar to SCAAT Welch and Bishop (1997).
@@ 3.2.2 Reasoning over states
Spatial relationship reasoning evaluates a model’s ability to infer qualitative spatial relationships between geometric objects. Spatial predicates defined in ArcGIS Johnston et al. (2001); Scott and Janikas (2009) lead to 35 possible combinations of geometry pairs and spatial relations (Appendix A.5). For each instance, a pair of geometric objects is sampled and a specific relationship is generated using ShapelyGillies and others (2007).
Temporal interval reasoning tests a model’s ability with Allen’s interval algebra Allen (1983). We generate tasks by simulating two random intervals, labeling their temporal relationship, and creating a balanced negative counterpart.
@@ 3.2.3 Knowledge-aware reasoning
Using ArcGIS Johnston et al. (2001); Scott and Janikas (2009) and OpenStreetMap Haklay and Weber (2008), we produce the following challenges that require world knowledge consideration: (1) Landmark direction evaluates a model’s ability to infer directional relationships between real-world landmarks. (2) Landmark proximity tests a model’s spatial reasoning in urban contexts by reasoning the presence of amenities near landmarks (e.g., Is there a police station within 500 meters of Central Park?). (3) Route planning tests whether models can connect textual travel descriptions to real-world spatial layouts. The model must decide if a given route passes by a specified landmark. (4) For ETA calculation, the model is asked to determine, given a departure time and average speed, if arrival before a specified deadline is possible based only on textual descriptions. (5) Route segment duration test the model on whether a user could plausibly be present along the segment during a given time window.
Additionally, we resort to simulating user trajectories for controlled ground-truth user intent to address data scarcity and privacy concerns in real-world mobility datasets. (6) POI prediction simulates human mobility using a finite state machine (FSM)Fernández-Caballero et al. (2012); Kerr et al. (2011), where each state represents a point of interest (POI). We generate a synthetic seven-day trajectory at hourly intervals. To enhance realism, we integrate contextual factors, such as weather, time, traffic, social events, etc., which are also expressed in natural language, into the models. The model must predict the next likely POI, requiring reasoning over spatiotemporal traces and language context (Figure 3 (b)). (7) Similarly, intent prediction extends POI prediction by requiring the model to infer the user’s underlying intent behind each visit. The model must predicts activity category (e.g., going home, watching a movie).
@@ 3.3 Baseline generation
We establish first-principle baselines for Tier 1 tasks using theoretically grounded methods for localization, tracking, and field variable prediction (details in Appendix A.6). For Tier 2 and Tier 3 tasks, however, many problems are solvable deterministically or through oracle-level services. For instance, spatial reasoning tasks such as point-in-polygon, intersection, or containment can be exactly solved using spatial libraries like Shapely Gillies and others (2007), while world-knowledge-based queries (e.g., nearest landmark, directional relation) can be precisely answered by GIS services such as ArcGIS Johnston et al. (2001). As these tools achieve near-perfect accuracy, including them as baselines would render comparisons with LLMs/LRMs uninformative. Instead, we report a random-guess baseline (0.5 accuracy under balanced labels) to reflect the challenge of learning such reasoning capabilities from language and context without deterministic priors.
@@ 4.1 Experiment setup
We conduct evaluations on both versions of the STARK benchmark. Specifically, we evaluated STARK-S (partial, 1.3k samples) and STARK-L (full, 14k samples), and we report results on STARK-S in the main paper and refer readers to the Appendix (Table 10 and 8) for detailed results on STARK-L and STARK-S. All system prompts are provided in Table 3. Cost information is shown in Table 4 to support dataset selection.
For all tasks except field variable prediction, we report the Root Mean Squared Error (RMSE) against the ground truth (for both open-ended calculation and true or false questions). For field variable prediction, where outputs vary in scale, we instead report the Root Mean Squared Percentage Error (RMSPE), because of differences in data magnitude across datasets. To ensure robustness to extreme outliers, we compute all metrics using the trimmed values unless stated otherwise, i.e., excluding the top and bottom 10% of the distribution.
For evaluating the models using a code interpreter (CI), we execute the code in a sandbox environment and report the relative change in error compared to direct answering (DA). This is calculated as the percentage change in RMSE and RMSPE due to using CI (lower values indicate better performance). Code execution was supported using scipy, numpy, shapely, and filterpy libraries Virtanen et al. (2020); Harris et al. (2020); Gillies and others (2007); Labbe (2014).